OpenStudy (anonymous):

x^3+8=0 What are the solutions and how do you get them? x^3+8=0 What are the solutions and how do you get them? @Mathematics

OpenStudy (anonymous):

x^3=-8 x=-2

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

are you looking for complex solutions as well? or just real?

OpenStudy (anonymous):

complex too

OpenStudy (anonymous):

then you should factor as a sum of two cubes:\[x^3+8=(x+2)(x^2+2x+4)\]One solution is the -2, the other are the two complex solutions of:\[x^2+2x+4=0\]which can be obtained from the quadratic formula.

OpenStudy (anonymous):

awesome thanks